# Massive Black Hole Seeds

### v1

@@ 9 Outlook
The next generation of simulations are likely to push this mass resolution limit even further slowing encroaching into the regime of the light seed (simulations by Bhowmick et al. (2024) already incorporate some of this thinking, albeit with optimistic dynamics as BHs are pinned to halo centres, increasing the chances of growth by both accretion and mergers). By including a self-consistent treatment of both light and heavy seeds and their dynamics, the growth and relative number densities of each seed population can in principle be determined. However, expectations must be tempered. For the foreseeable future a subgrid model for both light and heavy seed formation will exist particularly as the formation pathways for heavy seeds via collisional runaway (e.g. Arca Sedda et al., 2023; Rantala, Naab & Lahén, 2024) and via SMSs remain unconverged in detailed high resolution simulations (e.g. Prole et al., 2023).
@@
From the observational point of view, JWST has detected a new population of AGN, selected spectroscopically via broad (Maiolino et al., 2024b; Matthee et al., 2023) or narrow emission lines (Scholtz et al., 2023), or photometrically via color-color selection and then followed up spectroscopically to identify broad emission lines (Greene et al., 2024). This population has a number density above the extrapolation of the previously-known bright end of the AGN luminosity function: this is because these AGN have little or no X-ray emission (Furtak et al., 2023; Yue et al., 2024b) and they are UV-faint with respect to the host galaxies (Matthee et al., 2023; Scholtz et al., 2023). The physical reason for this is still unclear, but previous selections were based on X-ray and restframe-UV, and therefore this population was missed. Interestingly, reanalysis of HST images using variability has uncovered a similarly sized population that is UV bright (Hayes et al., 2024) and some of the high-z AGN are X-ray bright (Bogdán et al., 2023). However, as a population, the JWST AGN, or perhaps candidates, do not show X-ray emission (Yue et al., 2024a; Maiolino et al., 2024a; Ananna et al., 2024). This lack of X-rays has led to different hypotheses: the sources are not AGN, they’re AGN but the masses of MBHs have been overestimated, or they are intrinsically X-ray weak, or they are surrounded by dust-free Compton-thick gas.

### v2

@@ 9 Outlook
Semi Analytic Models (SAMs) rely on a backbone of dark matter merger trees to model galaxy and MBH evolution over long timescales. Their relative simplicity, compared to hydrodynamic simulations make them ideal for rapidly traversing the large parameter space of galaxy and black hole formation (e.g. Schneider et al., 2006; Dayal et al., 2019; Trinca et al., 2022). However, this power comes at the cost of small scale details. SAMs inherently do not have spatially-resolved structures and the detailed physical modelling capacity necessary to capture emergent physics like turbulence, detailed ISM physics as well as the likely episodic nature of MBH accretion.
In Figure 6 we show a schematic outlining the range of spatial and temporal scales that must be bridged by our models. No one simulation suite, technique or code can cross the domains and multiple suites and techniques (SAMs + hydrodynamics) must be used. To model MBH evolution different models utilise different seeding and accretion prescriptions - usually driven by the numerical resources available to them at a given scale. For example the large scale simulations (top right) must seed galaxies with relatively large black holes which represent black holes not necessarily at their seed masses but at some time after their physical formation.
SAMs based on high-resolution parent DM halo trees can in principle seed MBHs in arbitrarily low-mass galaxies/halos. However, SAMs have to seed BHs based on global quantities rather than local, spatially-resolved properties. Common to both the large scale hydrodynamic simulations and the SAMs is a difficulty in resolving the dynamics and detailed accretion physics of the MBHs. As we negotiate the parameter space of the schematic moving from top right to bottom left the models can becomes increasingly sophisticated in how, and at what mass, black holes are seeded. At smaller scales and higher resolution emergent physics gives rise to a better handle on the physics thought necessary for MBH formation (be they light or heavy seeds). Conversely, these simulations are unable, to follow over long timescale, the evolution of the seeds.
@@
The next generation of simulations are likely to push this mass resolution limit even further slowing encroaching into the regime of the light seed (simulations by Bhowmick et al. (2024a, b) already incorporate some of this thinking, albeit with optimistic dynamics as BHs are pinned to halo centres, increasing the chances of growth by both accretion and mergers). By including a self-consistent treatment of both light and heavy seeds and their dynamics, the growth and relative number densities of each seed population can in principle be determined. However, expectations must be tempered. For the foreseeable future a subgrid model for both light and heavy seed formation will exist particularly as the formation pathways for heavy seeds via collisional runaway (e.g. Arca Sedda et al., 2023; Rantala, Naab & Lahén, 2024) and via SMSs remain unconverged in detailed high resolution simulations (e.g. Prole et al., 2024).
@@
From the observational point of view, JWST has detected a new population of AGN, selected spectroscopically via broad (Maiolino et al., 2024b; Matthee et al., 2024) or narrow emission lines (Scholtz et al., 2023), or photometrically via color-color selection and then followed up spectroscopically to identify broad emission lines (Greene et al., 2024). This population has a number density above the extrapolation of the previously-known bright end of the AGN luminosity function: this is because these AGN have little or no X-ray emission (Furtak et al., 2024; Yue et al., 2024b) and they are UV-faint with respect to the host galaxies (Matthee et al., 2024; Scholtz et al., 2023). The physical reason for this is still unclear, but previous selections were based on X-ray and restframe-UV, and therefore this population was missed. Interestingly, reanalysis of HST images using variability has uncovered a similarly sized population that is UV bright (Hayes et al., 2024) and some of the high-z AGN are X-ray bright (Bogdán et al., 2023). However, as a population, the JWST AGN, or perhaps candidates, do not show X-ray emission (Yue et al., 2024a; Maiolino et al., 2024a; Ananna et al., 2024). This lack of X-rays has led to different hypotheses: the sources are not AGN, they’re AGN but the masses of MBHs have been overestimated, or they are intrinsically X-ray weak, or they are surrounded by dust-free Compton-thick gas.